# Assessing Small Language Models for Code Generation: An Empirical Study with Benchmarks

### v3

@@ Abstract
Small Language Models (SLMs) are lightweight variants of Large Language Models (LLMs), designed for simplicity, efficiency, and ease of use in resource-constrained environments [1]. Unlike LLMs, which contain hundreds of billions or even trillions of parameters, SLMs typically range from 1 million to 10 billion parameters [2]. Despite being significantly smaller, SLMs have shown effective performance on domain-specific tasks such as code generation [3]. This has led to growing interest in SLMs, as they offer a good balance between performance and efficiency [4]. With lower memory usage, faster inference, and easier deployment, SLMs can be adopted as a practical and sustainable alternative [5]. Therefore, to support SLMs further development and successful adoption in the real world, it is necessary to assess their performance.
Although SLMs are getting more attention, our understanding of their performance across code generation remains limited. Most studies so far have focused on LLMs such as Codex [6] and CodeLlama [7], while leaving smaller models relatively underexplored. Moreover, the few existing results often lack consistency in terms of benchmarks used (e.g., HumanEval, MBPP), programming language diversity (e.g., Python, Java, C++), and reporting of computational efficiency (e.g., memory usage, inference speed). This makes it difficult to compare results or draw meaningful conclusions about their effectiveness. Without clear testing, it is challenging to develop a complete and reliable understanding of SLMs and their real-world capabilities.
Despite recent progress in the development of SLMs, the lack of comprehensive evaluations [8] leaves many open questions about their actual effectiveness in practice. This gap motivates our study. To address this, we conduct a systematic evaluation aimed at building a better understanding of how these models perform across a range of conditions that reflect real world use. Specifically, we focus on three key aspects: (i) performance of SLMs on code generation, (ii) how they balance computational efficiency and performance, and (iii) how they perform in different programming languages. By examining these dimensions, we aim to provide insights into the strengths, weaknesses, and practical trade-offs of SLMs.
@@
A systematic evaluation of 20 open-source SLMs on standardised code generation benchmarks.
Detailed measurements of selected SLMs in terms of computational efficiency, including GPU memory consumption and inference speed.
Evaluation of the language-specific performance of selected SLMs to understand the multilingual capabilities and limitations of small language models in code generation tasks.
Paper organisation: Section 2 outlines the research methodology, which describes the experimental design, including selected models, benchmarks, and evaluation metrics. Section 3 presents the experimental findings, while Section 4 discusses their implications. Section 5 addresses potential threats to the validity of the results of this study. Section 6 reviews related work. Section 7 summarises the contributions and concludes the paper.
@@ 2.1 Research Design Overview
This study adopts an empirical research methodology using benchmark-based evaluation to assess the performance of SLMs in code generation tasks. Our study is structured around three phases: i) selecting SLMs and benchmarks, ii) evaluating the models across multiple benchmarks using a controlled experimental setup, and iii) performing quantitative data analysis. This research design addresses our three core research questions, focusing on SLMs i) performance, ii) computational efficiency, and iii) multilingual capabilities. It involves a systematic evaluation of selected SLMs through 5 benchmarks to ensure reproducible comparisons. Figure 1 provides a high-level overview of the research methodology.
@@ 2.2 SLMs Selection
We systematically selected 20 SLMs for this study using the following criteria. i) We selected SLMs released between 2022 and 2024 to ensure the study captures the latest advancements in the field. ii) We prioritised SLMs with high community engagement, using metrics such as Huggingface downloads and GitHub stars to reflect popularity. iii) We included only SLMs with open-source licenses (e.g., MIT, Apache 2.0) to support accessibility. iv) We restricted our selection to SLMs with up to 10 billion parameters[2] and categorised them into three groups, i.e., a) SLMs parameter size 1.5B, b) SLMs parameter size 1.5B to 3B, and c) SLMs parameter size 3B to 10B. v) We considered only SLMs that follow a decoder-only transformer architecture, which is widely used in autoregressive code generation tasks due to its effectiveness and simplicity [9]. Importantly, all our selected models are specifically designed for code generation tasks, ensuring relevance to the study’s focus on code generation performance. Table 1 provides an overview of the selected SLMs, including their specifications and metadata.
@@ 2.3 Benchmark Selection
We systematically selected five benchmarks, based on the following considerations: i) We prioritised benchmarks that are well-established in the research community and frequently used for evaluating LMs on code-related tasks [10, 11]. This ensures that our results are comparable with previous studies (e.g., Li et al. [12], Zhuo et al. [13]) and can be trusted by the broader research community. ii) We ensured diversity in evaluation tasks, including code synthesis, code summarisation, and code repair, to have a comprehensive assessment of SLMs capabilities. This variety reflects the multifaceted nature of software development, enabling us to assess how the SLMs handle different types of code generation challenges. iii) We selected benchmarks that support multiple programming languages, including Python, Java, JavaScript, PHP, Ruby, Go, and C++, to examine multilingual performance. iv) We included benchmarks that assess the SLMs outputs across various dimensions, such as functional correctness, and alignment with reference answers. These dimensions provide a holistic understanding of SLMs ability, balancing both performance and practical utility in real-world coding scenarios. Table 2 presents an overview of the five selected benchmarks, highlighting their key characteristics and relevance to the research questions.
2.4 Experimental Setup
2.4.1 Prompting Technique
We followed a zero-shot prompting strategy, where each model received only the raw benchmark input as a prompt, without supplementary examples or contextual hints. This design aims to assess the generalisation capabilities of each SLM and maintain consistency throughout the experiments.
2.4.2 Decoding Configuration
Consistent decoding parameters were maintained across all tasks. Adjustments were made only when certain benchmarks needed longer context lengths. To promote deterministic yet slightly different outputs, the temperature was set to 0.2, top-k sampling was disabled by setting it to 0, and nucleus sampling [14] was applied with top-p set to 0.95. The maximum generation length was fixed at 512 tokens for most tasks; however, for benchmarks demanding longer contexts, such as Mercury and HumanEvalPack (fixtests, fixdocs), this limit was extended to 2048 tokens. A default batch size of 10 was chosen to utilise all the available hardware resources. However, it was reduced to 5 for memory-intensive tasks to prevent out-of-memory errors. For CodeXGLUE, a batch size of 1 was adopted following its official evaluation guidelines.
| # | Model Name | Parameters | Architecture | License | Release Date | Company | Huggingface Downloads (Feb 2025) | GitHub Stars |
| Group 1: SLMs parameter size 1.5B | ||||||||
| 1 | PolyCoder | 0.4B | Decoder | MIT | Feb 2022 | Carnegie Mellon University | 273 | 1.8K |
| 2 | InCoder | 1.3B | Decoder | CC-BY-NC-4.0 | Apr 2022 | Facebook/Meta | 3060 | 3.2K |
| 3 | Phi 1 | 1.3B | Decoder | MIT | Jun 2023 | Microsoft | 7449 | N/A |
| 4 | DeepSeek-Coder | 1.3B | Decoder | DEEPSEEK-license | Nov 2023 | DeepSeek-AI | 123488 | 21K |
| 5 | OpenCodeInterpreter | 1.3B | Decoder | Apache 2.0 | Feb 2024 | Multimodal Art Projection (MAP) | 555 | 1.6K |
| 6 | Yi-Coder | 1.5B | Decoder | Apache 2.0 | Sep 2024 | 01.AI | 8913 | 7.8K |
| 7 | Qwen2.5-Coder | 1.5B | Decoder | Apache 2.0 | Sep 2024 | Alibaba Cloud | 11799 | 4.6K |
| Group 2: SLMs parameter size 1.5B to 3B | ||||||||
| 8 | CodeGemma | 2.0B | Decoder | Gemma | Apr 2024 | 37008 | N/A | |
| 9 | PolyCoder | 2.7B | Decoder | MIT | Feb 2022 | Carnegie Mellon University | 330 | 1.8K |
| 10 | Stable Code | 3.0B | Decoder | Apache 2.0 | Jan 2024 | Stability AI | 175 | 125 |
| 11 | StarCoder2 | 3.0B | Decoder | BigCode OpenRAIL-M | Feb 2024 | BigCode | 1468288 | 1.9K |
| 12 | Qwen2.5-Coder | 3.0B | Decoder | Qwen Research | Sep 2024 | Alibaba Cloud | 13540 | 4.6K |
| Group 3: SLMs parameter size 3B to 10B | ||||||||
| 13 | InCoder | 6.7B | Decoder | CC-BY-NC-4.0 | Apr 2022 | Facebook/Meta | 640 | 3.0K |
| 14 | DeepSeek-Coder | 6.7B | Decoder | DEEPSEEK-license | Nov 2023 | DeepSeek-AI | 28065 | 21K |
| 15 | OpenCodeInterpreter | 6.7B | Decoder | Apache 2.0 | Feb 2024 | Multimodal Art Projection (MAP) | 136 | 1.5K |
| 16 | CodeLlama | 7.0B | Decoder | Llama2 | May 2024 | Meta | 180516 | 16.2K |
| 17 | StarCoder2 | 7.0B | Decoder | BigCode OpenRAIL-M | Feb 2024 | BigCode | 125386 | 2.0K |
| 18 | CodeGemma | 7.0B | Decoder | Gemma | Apr 2024 | 42304 | N/A | |
| 19 | Qwen2.5-Coder | 7.0B | Decoder | Apache 2.0 | Sep 2024 | Alibaba Cloud | 64152 | 4.6K |
| 20 | Yi-Coder | 9.0B | Decoder | Apache 2.0 | Sep 2024 | 01.AI | 49711 | 7.8K |
@@ 2.4.4 Tooling and Automation
To ensure the overall evaluation process was smooth and efficient, we utilised the BigCode Evaluation Harness 111https://github.com/bigcode-project/bigcode-evaluation-harness, an open-source benchmarking framework that uses Huggingface Transformers and Accelerate libraries. This harness automates prompt formatting, response generation, and metric computation, making it a suitable choice for evaluating a large set of SLMs uniformly.
@@ 2.5 Metrics and Data Collection
is the number of correct (fully passing) samples,
@@
Peak memory usage, measured in megabytes, to assess resource demands.
@@
For RQ3, which explores language-specific model performance, we analysed benchmark results by programming language. HumanEvalPack and CodeXGLUE contain tasks written in multiple languages. For each language in HumanEvalPack, we computed the same pass@k and in CodeXGLUE, we calculated the BLEU score. BLEU is a precision-based metric commonly used in natural language generation to measure the overlap between generated and reference sequences. In our context, it quantifies how closely the generated code documentation matches the human-written descriptions, making it suitable for evaluating code-to-text tasks across different languages. This allowed us to examine whether models perform consistently across languages or exhibit biases.

### v4

@@ Abstract
Small Language Models (SLMs) are lightweight variants of Large Language Models (LLMs), designed for efficiency, and ease of use in resource-constrained environments [xu2024survey]. Unlike LLMs, which contain hundreds of billions or even trillions of parameters, SLMs typically range from 1 million to 10 billion parameters [Wang2024ACS]. Despite being significantly smaller, SLMs have shown effective performance on domain-specific tasks such as code generation [Subramanian2025survey]. This has led to growing interest in SLMs, as they offer a good balance between performance and efficiency [Nguyen2024survey]. With lower memory usage, faster inference, and easier deployment, SLMs can be adopted as a practical and sustainable alternative [schick-schutze-2021-just]. Therefore, to support SLMs further development and successful adoption in the real world, it is necessary to assess their performance.
Although SLMs are getting more attention, our understanding of their performance across code generation remains limited. Most studies so far have focused on LLMs such as Codex [Chen2021EvaluatingLL] and CodeLlama [CodeLlama], while leaving smaller models relatively underexplored. Moreover, the few existing results often lack consistency in terms of benchmarks used (e.g., HumanEval, MBPP), programming language diversity (e.g., Python, Java, C++), and reporting of computational efficiency (e.g., memory usage, inference speed). This makes it difficult to compare results or draw meaningful conclusions about their effectiveness. Without clear testing, it is challenging to develop a complete and reliable understanding of SLMs and their real-world capabilities.
Despite recent progress in the development of SLMs, the lack of comprehensive evaluations leaves many open questions about their actual effectiveness in practice [Lu2024SmallLM]. Existing studies mostly focus on general natural-language processing tasks [zellers2019hellaswag] [hendrycks2020measuring] [lewkowycz2022solving]. As a result, the effectiveness of SLMs for specialised software engineering tasks, such as code generation, remains insufficiently understood. Besides, the absence of cross-programming-language evaluations also limits understanding of whether SLM performance and efficiency vary across different programming languages. Moreover, prior work does not systematically examine task-specific trade-offs between code-generation quality and computational efficiency, including latency and memory usage, which are critical for practical deployment [Lepagnol2024SmallLM]. These gap motivates our study. To address these, we conduct a systematic evaluation aimed at building a better understanding of how these models perform across a range of conditions for practical usage. Specifically, we focus on three key aspects: (i) performance of SLMs on code generation, (ii) how they balance computational efficiency and performance, and (iii) how they perform in different programming languages. By examining these dimensions, we aim to provide insights into the strengths, weaknesses, and practical trade-offs of SLMs.
@@
A systematic evaluation of 20 open-source SLMs (i.e., 0.4B–10B parameters) divided into three groups across five widely used benchmarks, HumanEval, MBPP, Mercury, HumanEvalPack, and CodeXGLUE—using a unified decoding and prompting setup.
Detailed performance measurements of selected SLMs, including functional correctness (pass@k), GPU memory consumption, and inference speed, to better understand the computational trade-offs inherent in these models.
Evaluation of language-specific performance across selected SLMs, assessing the multilingual capabilities and limitations of small language models in code generation tasks.
Our results show that larger SLMs from Group 3 (parameter size 3B to 10B) generally deliver the highest accuracy, but several smaller and mid-sized models from Groups 1 and 2 (parameter size 3B) perform competitively while requiring substantially fewer resources. VRAM usage increases sharply with model size, although inference times remain statistically similar. While multilingual performance varies across individual languages, these differences are not statistically significant within the language set included in our benchmarks. Overall, the findings indicate important trade-offs between accuracy and efficiency that practitioners must consider when selecting SLMs for practical deployment.
Paper Organization: Section 2 outlines the research methodology, which describes the experimental design, including selected models, benchmarks, and evaluation metrics. Section 3 presents the experimental findings, while Section 4 discusses their implications. Section 5 addresses potential threats to the validity of the results of this study. Section 6 reviews related work. Section 7 summarizes the contributions and concludes the paper.
@@ 2.1 Research Design Overview
This study adopts an empirical, benchmark-based methodology to evaluate SLM performance in code generation. The workflow consists of three phases: (i) selecting SLMs & Benchmarks, (ii) Experimental Setup, and (iii) Data Analysis. This design directly supports our three research questions on model performance, computational efficiency, and multilingual capability, and ensures reproducible comparisons across five benchmarks. Figure 1 provides an overview of this process. Phase 1 covers the selection of 20 open-source SLMs and five benchmarks based on release date, licence, architecture, community support, and language coverage. Phase 2 illustrates the unified experimental setup, including zero-shot prompting, decoding configurations, hardware platforms, and automated evaluation. Phase 3 summarises data extraction and analysis, including metric computation (pass@k, BLEU, VRAM, inference time) and statistical testing. Together, these phases form a consistent and structured evaluation pipeline.
@@ 2.2 SLMs Selection
We systematically selected 20 SLMs for this study using the following criteria. i) We selected SLMs released between 2022 and 2024 to ensure the study captures the latest advancements in the field. ii) We prioritised SLMs with high community engagement, using metrics such as Huggingface downloads and GitHub stars to reflect popularity. iii) We included only SLMs with open-source licenses (e.g., MIT, Apache 2.0) to support accessibility. iv) We restricted our selection to SLMs with up to 10 billion parameters[Wang2024ACS] and categorised them into three groups, i.e., a) SLMs parameter size 1.5B, b) SLMs parameter size 1.5B to 3B, and c) SLMs parameter size 3B to 10B. v) We considered only SLMs that follow a decoder-only transformer architecture, which is widely used in autoregressive code generation tasks due to its effectiveness and simplicity [svyatkovskiy2020intellicode]. Importantly, all our selected models are specifically designed for code generation tasks, ensuring relevance to the study’s focus on code generation performance. Table 1 provides an overview of the selected SLMs, including their specifications and metadata.
| Model_Parameter | Architecture | License | Release Date | Company | Huggingface Downloads (Feb 2025) | GitHub Stars (Feb 2025) |
| Group 1: SLMs parameter size 1.5B | ||||||
| PolyCoder_0.4B | Decoder | MIT | Feb 2022 | Carnegie Mellon University | 273 | 1.8K |
| InCoder_1.3B | Decoder | CC-BY-NC-4.0 | Apr 2022 | Facebook/Meta | 3060 | 3.2K |
| Phi1_1.3B | Decoder | MIT | Jun 2023 | Microsoft | 7449 | N/A |
| DeepSeek-Coder_1.3B | Decoder | DEEPSEEK-license | Nov 2023 | DeepSeek-AI | 123488 | 21K |
| OpenCodeInterpreter_1.3B | Decoder | Apache 2.0 | Feb 2024 | MAP | 555 | 1.6K |
| Yi-Coder_1.5B | Decoder | Apache 2.0 | Sep 2024 | 01.AI | 8913 | 7.8K |
| Qwen2.5-Coder_1.5B | Decoder | Apache 2.0 | Sep 2024 | Alibaba Cloud | 11799 | 4.6K |
| Group 2: SLMs parameter size 1.5B to 3B | ||||||
| CodeGemma_2.0B | Decoder | Gemma | Apr 2024 | 37008 | N/A | |
| PolyCoder_2.7B | Decoder | MIT | Feb 2022 | Carnegie Mellon University | 330 | 1.8K |
| Stable Code_3.0B | Decoder | Apache 2.0 | Jan 2024 | Stability AI | 175 | 125 |
| StarCoder2_3.0B | Decoder | BigCode OpenRAIL-M | Feb 2024 | BigCode | 1468288 | 1.9K |
| Qwen2.5-Coder_3.0B | Decoder | Qwen Research | Sep 2024 | Alibaba Cloud | 13540 | 4.6K |
| Group 3: SLMs parameter size 3B to 10B | ||||||
| InCoder_6.7B | Decoder | CC-BY-NC-4.0 | Apr 2022 | Facebook/Meta | 640 | 3.0K |
| DeepSeek-Coder_6.7B | Decoder | DEEPSEEK-license | Nov 2023 | DeepSeek-AI | 28065 | 21K |
| OpenCodeInterpreter_6.7B | Decoder | Apache 2.0 | Feb 2024 | MAP | 136 | 1.5K |
| CodeLlama_7.0B | Decoder | Llama2 | May 2024 | Meta | 180516 | 16.2K |
| StarCoder2_7.0B | Decoder | BigCode OpenRAIL-M | Feb 2024 | BigCode | 125386 | 2.0K |
| CodeGemma_7.0B | Decoder | Gemma | Apr 2024 | 42304 | N/A | |
| Qwen2.5-Coder_7.0B | Decoder | Apache 2.0 | Sep 2024 | Alibaba Cloud | 64152 | 4.6K |
| Yi-Coder_9.0B | Decoder | Apache 2.0 | Sep 2024 | 01.AI | 49711 | 7.8K |
@@ 2.3 Benchmark Selection
We systematically selected five benchmarks, based on the following considerations: i) We prioritised benchmarks that are well-established in the research community and frequently used for evaluating LMs on code-related tasks [10606356, he-etal-2024-ultraeval]. This ensures that our results are comparable with previous studies (e.g., Li et al. [li2023starcoder], Zhuo et al. [zhuo2024bigcodebench]) and can be trusted by the broader research community. ii) We ensured diversity in evaluation tasks, including code synthesis, code summarisation, and code repair, to have a comprehensive assessment of SLMs capabilities. This variety reflects the multifaceted nature of software development, enabling us to assess how the SLMs handle different types of code generation challenges. iii) We selected benchmarks that support multiple programming languages, including Python, Java, JavaScript, PHP, Ruby, Go, and C++, to examine multilingual performance. iv) We included benchmarks that assess the SLMs outputs across various dimensions, such as functional correctness, and alignment with reference answers. These dimensions provide a holistic understanding of SLMs ability, balancing both performance and practical utility in real-world coding scenarios. Table 2 presents an overview of the five selected benchmarks, highlighting their key characteristics and relevance to the research questions.
2.4 Experimental Setup
2.4.1 Prompting Technique
We adopted a zero-shot prompting strategy for all benchmarks, where each model received only the raw benchmark input without supplementary examples or contextual hints. Zero-shot prompting is the standard evaluation protocol in widely used code generation benchmarks such as HumanEval [Chen2021EvaluatingLL] and MBPP [austin2021program], and is also commonly used in recent evaluations of code models (e.g., [li2023starcoder]). Before finalizing this approach, we considered whether alternative prompting strategies (e.g., few-shot prompting) would be appropriate for this study. However, prior work suggests that few-shot or example-driven prompting is primarily enabled by instruction tuning [ouyang2022training], and that base models do not reliably benefit from demonstration-based prompts. All SLMs evaluated in this study are base models, so incorporating few-shot prompts would introduce methodological bias and reduce comparability across models. For this reason, we intentionally excluded few-shot prompting from the experimental design. We acknowledge, however, that this choice limits our ability to assess model behavior in interactive or example-driven workflows, a limitation further discussed in section 5.
2.4.2 Decoding Configuration
Consistent decoding parameters were maintained across all tasks (see Table 3). Adjustments were made only when certain benchmarks needed longer context lengths. To promote deterministic yet slightly different outputs, the temperature was set to 0.2, top-k sampling was disabled by setting it to 0, and nucleus sampling [holtzman2019curious] was applied with top-p set to 0.95. The maximum generation length was fixed at 512 tokens for most tasks; however, for benchmarks demanding longer contexts, such as Mercury and HumanEvalPack (fixtests, fixdocs), this limit was extended to 2048 tokens. A default batch size of 10 was chosen to utilise all the available hardware resources. However, it was reduced to 5 for memory-intensive tasks to prevent out-of-memory errors. For CodeXGLUE, a batch size of 1 was adopted following its official evaluation guidelines. When computing pass@k, for all benchmarks we generated 10 samples(n = 10) samples per task. Multiple samples are required because SLMs outputs are nondeterministic, and using a single sample would produce unstable estimates of functional correctness. The choice of n = 10 follows the standard configuration used in the official evaluation guidelines, ensuring comparability with prior work while keeping the computational cost manageable. To support full reproducibility, all scripts used for running experiments and analysis are publicly available at: https://github.com/GPT-Laboratory/Assessing-Small-Language-Models-SLMs-for-Code-Generation.
| Parameter | Value / Description |
| Temperature | 0.2 |
| Top-k | 0 (disabled) |
| Top-p | 0.95 |
| Max generation length | 512 tokens (default) 2048 tokens for Mercury and HumanEvalPack fixtests/fixdocs |
| Batch size | 10 (default across most tasks) 5 for memory-intensive tasks 1 for CodeXGLUE (official guidelines) |
| Number of samples (n) for pass@k | 10 |
@@ 2.4.4 Tooling and Automation
We used the BigCode Evaluation Harness111https://github.com/bigcode-project/bigcode-evaluation-harness, an open-source benchmarking framework built on top of Hugging Face Transformers and Accelerate to ensure a consistent, scalable, and reproducible evaluation workflow across all 20 SLMs. The framework provides a unified interface for running code-generation benchmarks and automates several parts of the evaluation pipeline. In our setup, it was responsible for loading models and tokenizers, formatting prompts according to each benchmark’s specification, executing model inference, and computing metrics such as pass@k and BLEU score [papineni2002bleu]. By standardizing these steps, the framework ensures that all models are evaluated under identical conditions.
For acceleration, the framework relies on the HuggingFace Accelerate backend, which manages device placement, memory allocation, and mixed-precision execution. All models in this study were executed on a single GPU using PyTorch with BF16 precision, which reduces memory usage while maintaining numerical stability during generation. Accelerate automatically handles casting, GPU memory safety checks, and synchronised execution without requiring manual optimisation. Importantly, we did not employ additional performance-enhancing frameworks such as vLLM, TensorRT, FlashAttention, quantisation, or custom fused kernels. This ensures that all models were evaluated using the same baseline acceleration settings, avoiding discrepancies that could arise from specialised optimisation techniques.
@@ 2.5 Metrics and Data Collection
is the number of correct samples,
@@
Peak memory usage, measured in gigabytes, to assess resource demands.
@@
For RQ3, which explores language-specific model performance, we analysed benchmark results by programming language. HumanEvalPack and CodeXGLUE contain tasks written in multiple languages. In HumanEvalPack, models are instructed to generate code, on the other hand, for CodeXGLUE, models are generating code documentation in natural language. For each tasks in each programming language in HumanEvalPack, we computed the same pass@k and in CodeXGLUE, we calculated the BLEU score. BLEU is a precision-based metric commonly used in natural language generation to measure the overlap between generated and reference sequences. In the multilingual evaluation, the big-code evaluation harness compute the BLEU score following the standard n-gram precision–based formulation with a brevity penalty. For a candidate generation and reference solution , the BLEU score is calculated as:
where denotes the modified n-gram precision, is the uniform weight for each n-gram order, and BP is the brevity penalty:
In our context, it quantifies how closely the generated code documentation matches the human-written descriptions, making it suitable for evaluating code-to-text tasks across different languages. This allowed us to examine whether models perform consistently across languages or exhibit biases.